You have just purchased a newly issued 90-day bank-accepted bill with a face value of $100,000. Given you paid exactly $99,000 for the bill, what was your annual nominal required rate of return on debt?



1
Expert's answer
2020-11-09T06:54:49-0500

"nominal rate of return on the bill (90days)= \frac {Face value-Cost of the bill} {Cost of the bill}"



Face value= $100,000

Cost of the bill= $99,000


"nominal rate of return on the bill (90 days) = \frac{100,000-99,000}{99,000}" = 0.01= 1%


convert to annual nominal rate of return ="1 \times 4" = 4%


Therefore the annual nominal rate of return = 4 %